OpenStudy (anonymous):

A circle has a radius of 8 inches. Find the area of a sector of the circle if the sector has an arc that measures 45°.

hero (hero):

calistar, have you considered applying the general formula I gave you?

OpenStudy (anonymous):

srry i still dnt understand=/

hero (hero):

What is it that you don't understand?

OpenStudy (anonymous):

correct me if im wrong, is it 8pi squared inches

hero (hero):

How did you get that?

hero (hero):

Show me your work.

OpenStudy (anonymous):

it was a educational guess

OpenStudy (2bornot2b):

\[\frac{[(\pi\times8^2)(8\times \pi/2)]}{(2\pi\times8)}\]

hero (hero):

Calistar, I'm only going to show you this once. If you don't get it, tell me what it is that you don't get. I'm going to explain it to you the best way that I can.

hero (hero):

In general, you are given a circle, a radius, and the angle, θ, that cuts across a given arc:  and you are essentially asked to find the length of arc AC. The general formula to find the length of arc AC is given by the following: \[\frac{\theta}{360°} = \frac{x}{2 \pi r}\], x = length of arc AC. Solving for x gives you the length of arc AC. In this particular case, you are given θ = 45°, and r = 8, and asked to find the arc length, x, of the portion of the circle that θ = 45° cuts across. It would look something like this:  Since we are given some of the components of the general formula such as θ = 45° and r = 8, we can substitute that into the formula and solve for x. Here's the setup: \[\frac{45}{360} = \frac{x}{2 \pi 8}\]. All you need to do at this point is to cross multiply and solve for x. Do you believe you can do this?